OpenStudy (anonymous):

cube root of 56 x to the fourteenth power.

OpenStudy (anonymous):

what are the directions to this? o.o is there any other info to this problem ? what topic / subject does this fall under?

OpenStudy (anonymous):

algebra 2; radicals

OpenStudy (agreene):

\[\large (\sqrt[4]{56x})^{14}\] is this the question?

OpenStudy (agreene):

make that 4 a 3, cubic... yeah

OpenStudy (anonymous):

yes

OpenStudy (agreene):

Well, note that this is the same as \[\large((56x)^{\frac14})^{\frac1{14}}\] try and remember the rules of exponents, if you get stuck lemme know.

OpenStudy (anonymous):

is the answer 2x^4 cube root 7x^2 ?

OpenStudy (phi):

\[\large((56x)^{\frac13})^{14} \]

OpenStudy (anonymous):

thats the answer ?

OpenStudy (phi):

No, just making sure I have the question right. I'd break it into \[ (56x)^{\frac{12}{3}} (56x)^{\frac23} \]

OpenStudy (phi):

first part gives us (56x)^4

OpenStudy (phi):

for the second part : (56x)^(2/3)= (7*2^3)^(2/3) = 7^(2/3) * 2^2 = 4* 49^(1/3)

OpenStudy (anonymous):

the question is ^3\[\sqrt{56x}\]^14

OpenStudy (phi):

Yes the problem is \[(\sqrt[3]{56x})^{14}\]

OpenStudy (phi):

which is the same as \[ (56x)^{\frac{14}{3}} \]

OpenStudy (phi):

\[4(56x)^4\sqrt[3]{49x^2}\]

OpenStudy (phi):

forgot the x^2

OpenStudy (phi):

Did I lose you?